OpenStudy (anonymous):

How do I complete the square? X^2+14x+_____

OpenStudy (anonymous):

x(x+14) x=0 x=-14 am i right ? sorry, if i made mistakes.

OpenStudy (anonymous):

Hmm, no. On the reveiw answer sheet she gave 14 • 1/2=(7)^2=49 I just don't remember how to work it...

OpenStudy (zehanz):

Remember that (a + b)² = a² + 2ab + b². If you see x² + 14x + ... as the result of this rule, it has to be (x + 7)², because (x + 7)² = x² + 14x + 49. 14x is 2*x*7, just as in the 2ab term. That is why you know it has to be x + 7. 7² = 49 is the last term, same as the b². So 49 completes the square. Hope this helps! ZeHanz.